It took Sara's mother 5 hours round trip to drive to the University and bring Sara back home for spring break. If the university is 137 1/2 miles from home, find her mothers average speed.
Answer by checkley77(12844)

(2*137.5)/5=275/5=55 mph. was her average speed.